OpenStudy (anonymous):

3x - 5y =9; x =-1,0,1

OpenStudy (anonymous):

Substitute ? i guess?

OpenStudy (osanseviero):

Ok, so the first step is to have y alone so you can put the value of x. (Like in the other question)

OpenStudy (anonymous):

so itll be 3x-5=9

OpenStudy (osanseviero):

No. You need to do arithmetic. Add, substact or divide so you can have the y alone

OpenStudy (osanseviero):

In the form of y = a +bx (a and b are numbers)

OpenStudy (osanseviero):

In : 3x - 5y =9 The first step would be to add 5y to each side

OpenStudy (anonymous):

3x - 5y =9; x =-1,0,1 is it X equals to the following? -1 , 0 , 1 is that it?

OpenStudy (anonymous):

if it is you can have the 3 answers for Y ? first as the other guy said make the Y alone . Amm LIke this >>>> y = (9 - 3x)/-5 <<<< Then substitute the X .